OpenStudy (anonymous):

h(t)=80+64t -16t. How do I solve for time

OpenStudy (matt101):

Do you know the quadratic formula?

OpenStudy (radar):

Simplify by combining similar terms getting h(t) = 48t + 80 Now use algebra: 48t = h(t)-80 t = (h(t)-80)/48 Obviously you need to know the value of h(t) to solve for t.

OpenStudy (anonymous):

Okay, that helped a whole lot. Thank you so much :)